OpenStudy (anonymous):

Use graphs and tables to find the limit and identify any vertical asymptotes of

OpenStudy (anonymous):

@zpupster

OpenStudy (zpupster):

the language as x approaches 3 from the left ( becasue of the minus sign) it can not reach 3 then it would be undefined so we look at values close to three but also left of three 2.9-3 = -.1 2.99-3 = -.01 2.999-3 = -.001 and so on we see that as we approach 3 we approach a negative number increasingly small or \[-\infty\]

OpenStudy (anonymous):

so x would equal 3 right? @zpupster sorry openstudy stopped working for a second

OpenStudy (anonymous):

-3*

OpenStudy (zpupster):

the language as x approaches 3 from the left 1/(x-3) approaches −∞

OpenStudy (anonymous):

wait im confused is x= -3 or x=3 ? I think it is -3 because it is approaching from the left

OpenStudy (anonymous):

@zpupster

OpenStudy (anonymous):

would it be written like A or B? @ganeshie8

OpenStudy (zpupster):

maybe a graph will help you see as we go to three we go to −∞

OpenStudy (anonymous):

so x=3 is correct?

OpenStudy (zpupster):

no −∞ is correct

OpenStudy (anonymous):

there are two parts to this question!

OpenStudy (anonymous):

i understand that, but x equals something

OpenStudy (anonymous):

and identify any vertical asymptotes

OpenStudy (anonymous):

the vertical asymptote would be x=3 right?

OpenStudy (anonymous):

@satellite73 please help!

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

The vertical asymptote is indeed 3..whatever makes the denominator = 0 will cause a vertical asymptote

OpenStudy (anonymous):

thank you!

OpenStudy (johnweldon1993):

Now to actually solve the limit, since we are approaching from the left side (or numbers lower than 3) we can just plug in 2.9 for 'x'...and then 2.99...and 2.999 so on and so forth, and see what the numbers approach